# High-probability Convergence Guarantees of Decentralized SGD

### v3

@@ Appendix H Deriving the Transient Time
Next, consider strongly convex costs. Recalling the expressions in (44) and (F), it follows that the full bound for strongly convex costs is of the form
$\mathcal{O}$ $\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{t_{0}^{3}}{n(T+t_{0})^{3}}$
@@
where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$. Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{1}{1-\lambda}\Big)$, so that the bound in (H) becomes
$\quad\quad\mathcal{O}\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}$
$+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{8}(T+t_{0})^{8}}\bigg).$ (50)
It can be seen that, for any $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (H) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [33], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., Table II in [1]. As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.
Appendix I Achieving Linear Speed-up
@@
where it is easy to see that choosing $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$ leads to the desired linear speed-up. Works like [44, 45, 76] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity. As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (52) and achieve linear speed-up. As discussed in Appendix E, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
@@
Appendix J Extending to Time-varying Networks
@@
Appendix K On Dimension Dependence
As mentioned in Section 4 in the main body, our rates in Theorems 1 and 2 exhibit a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, with the dependence stemming from Lemma 2, where it is shown that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\big)$-sub-Gaussian. This is a consequence of working with random vectors, where we simultaneoulsy need to show a bound on the MGF of the inner product $\langle\overline{z}^{t},v\rangle$ for any $\mathcal{F}_{t}$-measurable vector $v\in\mathbb{R}^{d}$ (shown in point 2 of Lemma 2), as well as on the MGF of the norm-squared $\|\overline{z}^{t}\|^{2}$. While the inner product maintains some desirable properties, such as being zero-mean and linear (in the sense that $\langle\overline{z}^{t},v\rangle=\frac{1}{n}\sum_{i\in[n]}\langle z^{t}_{i},v\rangle$), this is not the case with the squared norm, which is neither zero-mean, nor linear. Therefore, trying to directly establish the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma}{\sqrt{n}}\big)$-sub-Gaussian, in the sense of condition (A4), fails to yield the desired result. In particular, recalling that $\sigma^{2}=\frac{1}{n}\sum_{i\in[n]}\sigma_{i}^{2}$, we then have
@@
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\prod_{i\in[n]}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/n}\stackrel{{\scriptstyle(iv)}}{{\leq}}\prod_{i\in[n]}\exp\bigg(\frac{1}{n}\bigg)=\exp(1),$
@@
which holds for any $n\in\mathbb{N}$, $a_{i}\in\mathbb{R}$ and $b_{i}>0$, in $(iii)$ we used Proposition 1 and the fact that noise is conditionally independent across users, while $(iv)$ follows from (A4). Therefore, a direct approach yields that $\overline{z}^{t}$ is $\sigma$-sub-Gaussian, failing to show the variance reduction benefit. To circumvent this issue, we use a different argument, namely that a zero-mean random vector $x\in\mathbb{R}^{d}$, which is $\sigma$-sub-Gaussian in the inner product sense, i.e., for any $v\in\mathbb{R}^{d}$
$\mathbb{E}\big[\exp\big(\langle x,v\rangle\big)\big]\leq\exp\bigg(\frac{\sigma^{2}\|v\|^{2}}{2}\bigg),$
is also $\sigma\sqrt{d}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$
$\mathbb{P}\big(\|x\|>\epsilon\big)\leq 2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma^{2}d}\bigg),$
see Lemma 1 in [27]. This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence. One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian. Using a similar argument to the one in Lemma 6 and Corollary 7 in [27], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$. This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2. Finally, we note that, in the regime $n\ll d$, we can simply use the above argument following Sedrakyan’s inequality, to conclude that $\overline{z}^{t}$ is $\sigma$-sub Gaussian, reducing the dependence on problem dimension $d$, at the cost of losing linear speed-up in the number of users.

### v5

@@ Appendix I Deriving the Transient Time
Next, consider strongly convex costs. Recalling the expressions in (47) and (G), it follows that the full bound for strongly convex costs is of the form
$\mathcal{O}$ $\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{t_{0}^{3}}{(T+t_{0})^{3}}$
@@
where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$. Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{1}{1-\lambda}\Big)$, so that the bound in (I) becomes
$\quad\quad\mathcal{O}\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}$
$\quad\quad+\frac{1}{(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{8}(T+t_{0})^{8}}\bigg).$ (53)
It can be seen that for $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{\sqrt{n}}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (I) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$. Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [35], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., Table II in [1]. As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.
Appendix J Achieving Linear Speed-up
@@
where it is easy to see that choosing $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$ leads to the desired linear speed-up. Works like [47, 48, 83] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity. [^12] We note that [48] consider the problem of online noncooperative games and show their HP convergence guarantees in terms of dynamic regret $R_{i}(t)$, for each individual $i\in[n]$. While this is different from the metrics considered in our work (i.e., user-averaged gradient norm/optimality gap), the analysis in [48] suffers from the same shortcomings as the decentralized works, i.e., no result showing variance reduction in the HP sense is provided, making linear speed-up impossible to achieve even if an average of $R_{i}(t)$’s is considered. As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (55) and achieve linear speed-up. As discussed in Appendix F, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
@@
Appendix K Extending to Time-varying Networks
@@
Appendix L On Sub-Gaussian Vectors and Dimension Dependence
As mentioned in Section 4 in the main body, our rates in Theorems 1 and 2 exhibit a mild dependence on the problem dimension, of order $\log(d)$, with the dependence stemming from Lemma 2. This is a consequence of working with random vectors, where we simultaneoulsy need to show a bound on the MGF of the inner product $\langle\overline{z}^{t},v\rangle$ for any $\mathcal{F}_{t}$-measurable vector $v\in\mathbb{R}^{d}$ (shown in point 2 of Lemma 2), as well as on the MGF of the squared norm $\|\overline{z}^{t}\|^{2}$ (shown in point 3 of Lemma 2). While the inner product maintains some desirable properties, such as being zero-mean and linear (in the sense that $\langle\overline{z}^{t},v\rangle=\frac{1}{n}\sum_{i\in[n]}\langle z^{t}_{i},v\rangle$), this is not the case with the squared norm, which is neither zero-mean, nor linear. Therefore, trying to directly establish the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma}{\sqrt{n}}\big)$-sub-Gaussian, in the sense of condition (A4), fails to yield the desired result. In particular, recalling that $\sigma^{2}=\frac{1}{n}\sum_{i\in[n]}\sigma_{i}^{2}$, we then have
@@
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\prod_{i\in[n]}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/n}\stackrel{{\scriptstyle(iv)}}{{\leq}}\prod_{i\in[n]}\exp\bigg(\frac{1}{n}\bigg)=\exp(1),$ (59)
@@
which holds for any $n\in\mathbb{N}$, $a_{i}\in\mathbb{R}$ and $b_{i}>0$, in $(iii)$ we used Proposition 1 and the fact that noise is conditionally independent across users, while $(iv)$ follows from (A4). Therefore, a direct approach yields that $\overline{z}^{t}$ is $\sigma$-sub-Gaussian, failing to show the variance reduction benefit. To circumvent this issue, we use a different argument, namely that condition (A4) is equivalent to norm-sub-Gausiannity, i.e., that for any $\epsilon>0$, we have
$\mathbb{P}\big(\|z^{t}_{i}\|^{2}>\epsilon\>|\>\mathcal{F}_{t}\big)\leq 2\exp\bigg(\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg),$
which, combined with Proposition 9 (i.e., Corollary 7 in [29]), implies that
$\mathbb{P}\big(\|\overline{z}^{t}\|^{2}>\epsilon\big)\leq 2d\exp\bigg(\frac{n\epsilon}{3\sigma^{2}}\bigg).$
This result can be further leveraged to show the desired bound on the MGF of $\overline{z}^{t}$, while introducing the mild $\log(d)$ dependence, see the proof of Lemma 2 for details. It can be noted that, when there is no noise, i.e., $\sigma=0$, the dimension dependence vanishes, highlighting that this dependence stems from the noise. As mentioned in [29], while mild, it is not clear whether this dependence on $d$ can be completely removed.
We now highlight three important facts. First, if a different definition of sub-Gaussian vectors is used, namely the one via inner products, which states that, for any $\mathcal{F}_{t}$-measurable vector $v$, we have
$\mathbb{E}\big[\exp\big(\langle v,z^{t}_{i}\rangle\big)\>|\>\mathcal{F}_{t}\big]\leq\exp\bigg(\frac{\sigma_{i}^{2}\|v\|^{2}}{2}\bigg),$
this would inevitably lead to a worse dependence on the problem dimension, as in order to bound the MGF of the squared norm of the average noise, i.e., $\|\overline{z}^{t}\|^{2}$, one would need to use a similar argument as the one for Lemma 1 in [29], using covers of the unit sphere and picking up a factor of $d$ in the final rate, which is much worse than our $\log(d)$ dependence. Second, while the notion of sub-Gaussian random vectors used in (A4) is equivalent to the notion of norm-sub-Gaussian vectors from [29] and stronger than the notion of sub-Gaussian vectors defined via inner products above, it is in fact the standard notion of sub-Gaussianity used in almost all HP results, in both centralized and decentralized settings, see [58, 20, 40, 42, 44, 47, 48, 83] and references therein. Finally, we note that in the extreme regime $n\ll\log(d)$, we can simply use the argument in (L), to conclude that $\overline{z}^{t}$ is $\sigma$-sub-Gaussian, completely removing the dependence on dimension $d$, at the cost of losing linear speed-up in the number of users.